OpenStudy (anonymous):

Rewrite the following equation in the center-radius form of the equation of a circle. x^2 + 18x + y^2 - 18y = -146

OpenStudy (jdoe0001):

http://www.mathwarehouse.com/geometry/circle/images/equation-of-circle/general-formula-equation-of-circle.png do you know what a "perfect square trinomial" is?

OpenStudy (ranga):

Complete the squares of (x^2 + 18x) and (y^2 - 18y) separately first.

OpenStudy (jdoe0001):

ditto

OpenStudy (ranga):

I will do the first part and you can apply the same method. x^2 + 18x Divide the coefficient of the x term by 2: +18 / 2 = +9. This is the number that will go inside the square. Then you should subtract the square of this number: x^2 + 18x = (x + 9)^2 - 9^2 = (x + 9)^2 - 81 You try completing the square for y^2 - 18y